OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 2y = -6 3x - 2y = 11

OpenStudy (anonymous):

(-1, -7) (1, -4) (2, -1) (3, -2)

OpenStudy (jdoe0001):

so, graph them both, where they MEET, or TOUCH each other, that's the solution at THAT POINT, "x" and "y" for the top and bottom equations are the exact same value

OpenStudy (anonymous):

i dont need a graphing calculator

OpenStudy (jdoe0001):

well, ok, then don't use one :)

OpenStudy (anonymous):

ok im confused a little

OpenStudy (jdoe0001):

well, on what part?

OpenStudy (anonymous):

how to solve 2x+2y=-6

OpenStudy (jdoe0001):

well, you're suggested to use a graph, either on paper or electronic

OpenStudy (jdoe0001):

graph them both, where they MEET, or TOUCH each other, that's the solution at THAT POINT, "x" and "y" for the top and bottom equations are the exact same value

OpenStudy (anonymous):

i dont have electonic

OpenStudy (jdoe0001):

well, lemme see on that, you're sitting at a computer a computer is an electronic, in fact really just a large calculator so :/

OpenStudy (anonymous):

oh lol

OpenStudy (jdoe0001):

well, I guess to graph them first, you'd need to solve for "y" each so let's see what we get y = (-6-2x)/2 ==> y = -x-3 (3x-11)/2 =y ==> y = 3x/2 - 11/2

OpenStudy (jdoe0001):

so, those 2 equations

OpenStudy (anonymous):

i cant find it

OpenStudy (anonymous):

1,-4

OpenStudy (anonymous):

is that right

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

what platform are you on? macOSX, *nix, windows?

OpenStudy (anonymous):

windows 8

OpenStudy (jdoe0001):

ok, @Courage. http://graphcalc.com/download.shtml a graphic calculator